The electric potential energy stored in the capacitor of a defibrillator is 100 J, and the capacitance is 170 μF. What is the potential difference across the capacitor plates?
1
Expert's answer
2020-09-14T10:23:45-0400

"W=\frac{CU^2}{2}\to U=\sqrt\frac{2W}{C}=\sqrt\frac{2\cdot 100}{170\cdot10^{-6}}\approx1085(V)"